### v1

# Homogenized C. elegans Neural Activity and Connectivity Data

Abstract
There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C. elegans). This is particularly interesting as this model system provides a path to bridge the gap between structure and function, from nervous system connectivity to physiology. However, the many existing physiology datasets, both recording and stimulation, as well as connectome datasets, are in distinct formats, requiring extra processing steps before modeling or other analysis can commence. Here we present a homogenized dataset of neural activity, including during stimulation, compiled from 11 neuroimaging experiments and from 10 connectome reconstructions. The physiology datasets, collected under varying experimental protocols, all measure neural activity via calcium fluorescence in labeled subsets of the worm’s 300 neurons. Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling frequency. The resulting dataset includes neural recordings from approximately 900 worms and 250 uniquely labeled neurons. The connectome datasets, collected from electron microscopy (EM) reconstructions, all contain the entire nervous system of the worm, preprocessed into a graph of connections across the neurons. Using our collection of datasets is facilitated through easy data sharing on HuggingFace. We believe that our joint dataset of physiology and connectivity will facilitate modeling, for example in terms of recurrent neural network or transformer architectures, making it easier to check how well models generalize across animals and labs.
1 Introduction
Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience and machine learning. While neural dynamics in the brain can often be complex and difficult to measure in larger organisms, smaller systems like C. elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure [24, 3]. C. elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [25]. This makes it an ideal organism for studying how neural dynamics arise from connectivity and how this relationship could inform both neuroscience and AI models of neural function.
With its transparent body and well-characterized neural architecture, C. elegans provides an excellent model for whole-brain functional imaging using calcium fluorescence sensors [26]. Calcium indicators act as proxies for voltage activity, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6]. Additionally, since the full connectome of C. elegans has been mapped via electron microscopy [24, 25], combining these two datasets—neural dynamics and connectomic structure—allows for a deep investigation of how network connectivity influences neural function. Unification of such datasets can also contribute to the broader goal of understanding how neural activity propagates through networks, ultimately informing both the development of AI-based neural systems and providing insights into how these mechanisms generalize to more complex organisms [8].
However, the existing datasets from calcium imaging studies of C. elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates [9, 10]. This makes it challenging to perform cross-study comparisons or build unified models.
To address these challenges, we have created two datasets, the first combining neural dynamics from 11 neuroimaging studies and the second combining neural connectivity from 10 connectome studies of C. elegans. This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data [5]. The primary goal of this dataset is to facilitate the training of foundation models of neural dynamics, which can be used to study how neural function arises from structure. The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel [15]. It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function [3]. We are releasing these datasets as open-source resources on the HuggingFace platform to ensure broad accessibility to researchers in both neuroscience and machine learning. By unifying these datasets and making them accessible, we hope to encourage the development of computational models that can capture the structure-function relationships in small nervous systems.
2 Methods
2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that supports the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions and files) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data in a standardized format.
The main classes in the repository are NeuralBasePreprocessor and ConnectomeBasePreprocessor. The former manages processing of neural activity data as measured by calcium fluorescence imaging [23]. The latter manages the processing of synaptic connectivity data as measured by anatomical counts from electron microscopy [24].
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The former handles the extraction and processing of the raw calcium data from each of our 12 neural activity source datasets into standardized neural activity data that are saved as compressed files. The latter processes the raw connectome data from connectome source files into compressed graph tensors that are saved as compressed in a data format amenable to downstream processing by graph neural network packages. There are 7 distinct connectome source files as shown in Table 5 but they are all derived from 3 primary source publications [24, 3, 25].
Therefore, our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
2.2 Calcium Fluorescence Data
We processed neural activity datasets from 11 open-source studies of C. elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [14, 5, 18, 26, 9, 15, 12, 10, 11, 13]. These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals. The number of worms and identified neurons varies across datasets as showb in Table 1). All animals were hermaphrodites at developmental stages from L4 to adulthood.
Table 1: Calcium Fluorescence Neural Activity Datasets Metadata. Metadata for calcium fluorescence neural activity datasets collected from various sources. Each dataset includes information about the experimental conditions, the number of worms, the average number of labeled and recorded neurons, and the range of labeled neurons observed.
Source Dataset Name Database Link Data Files Num. Worms Mean Num. Neurons (labeled, recorded) Num. labeled neurons (min, max)
Kato2015 [9] osf.io/2395t/ WT_Stim.mat 12 (42, 127) (31, 51)
Nichols2017 [12] osf.io/kbf38/ let.mat 44 (34, 108) (23, 43)
Skora2018 [15] osf.io/za3gt/ WT_.mat 12 (46, 129) (39, 55)
Kaplan2020 [9] osf.io/9nfhz/ Neuron2019_Data_ *.mat 19 (36, 114) (23, 51)
Yemini2021 [26] zenodo.org/records/ 3906530 _Activity_OH.mat 49 (110, 125) (33, 179)
Uzel2022 [18] osf.io/3vkxn Uzel_WT.mat 6 (50, 138) (46, 58)
Dag2023 [13] tinyurl.com/github Dag2023 data/swf702_with_ id/*.h5 7 (100, 143) (87, 110)
Atanas2023 [5] wormwideweb.org YYYY-MM-DD-.json/h5 42 (88, 136) (64, 115)
Leifer2023 [14] tinyurl.com/drive Leifer2023 exported_data.tar.gz 108 (64, 68) (19, 98)
Lin2023 [11] tinyurl.comdropbox Lin2023 run_prfrd_data.mat 577 (8, 8) (1, 22)
Venkatachalam2024 [22] chemosensory-data.worm.world 2022_herm_.zip 22 (187, 187) (185, 189)
2.2.1 Neural Data Processing
The entire pipeline is implemented in the NeuralBasePreprocessor class, which handles dataset-specific preprocessing tasks, including file format loading, trace extraction, and metadata creation.
Table 2: Notation Used in the Neural Data Processing Pipeline. Description of the mathematical notation used throughout the neural data preprocessing pipeline. Variables represent worm-specific indices, neural activity data structures, temporal properties, neuron counts, and preprocessing parameters.
Symbol Description
$k$ Worm index variable within a single dataset
$\mathbf{X}^{(k)}$ Neural activity data for a single worm $k$
$T_{k}$ Number of time points of data recorded for a worm $k$
$D$ Number of neurons in C. elegans
$\mathbf{\mu},\mathbf{\sigma}$ Mean and standard deviation over time of neural activity data for each neuron
$\alpha$ Optional smoothing hyperparameter
$\mathbf{M}^{(k)}$ Binary mask generated for each worm using labeled neurons applied to neural activity data
The neural activity data is preprocessed using a custom pipeline (Figure 1) which follows these steps:
1. Data Download and Extraction. The raw data from each dataset is downloaded from its respective source and extracted. The function download_url() downloads the dataset, and extract_zip() decompresses the data into a local directory while preserving the original folder structure.
2. Custom Dataset Implementations. Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format. The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking.
We represent the neural activity data for a worm $k$ by a matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of time points and $D=300$ is the number of neurons in C. elegans. The neural activity at any time point $t$ is denoted as $\mathbf{X}^{(k)}[t]\in\mathbb{R}^{D}$, which represents the activity snapshot across the $D$ neurons.
3. Neuron Trace Normalization. The first preprocessing step is to normalize or z-score the raw (i.e. from the source datasets) neural activity across the neuron dimension. For each neuron trace $i$ in worm $k$, we compute:
$\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{raw}}-\mathbf{\mu}}{\mathbf{\sigma}}$
where $\mathbf{\mu},\mathbf{\sigma}\in\mathbb{R}^{D}$ are the mean and standard deviation over time for each neuron.
4. Smoothing. Optionally, the calcium traces can be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
$\mathbf{X}^{(k)}_{\text{smooth}}[t]=\alpha\mathbf{X}^{(k)}_{\text{norm}}[t]+(1-\alpha)\mathbf{X}^{(k)}_{\text{norm}}[t-1]$
where $\alpha$ is a smoothing hyperparameter set in preprocess.yaml. For our released neural dataset, smoothing was not applied ($\alpha=1$) as prior preprocessing had already been performed. Calcium fluorescence data inherently reflects low frequency neural dynamics. Smoothing, which acts further low pass over the raw data, is not needed.
5. Resampling. We resample the data to a fixed time interval $\Delta t$, using linear interpolation for datasets with different temporal resolutions. The new time index $t^{\prime}$ is calculated by:
$t^{\prime}=\left\lfloor\frac{t\cdot\Delta t_{\text{original}}}{\Delta t}\right\rfloor$
The resampled traces $\mathbf{X}_{\text{resampled}}^{(k)}$ are computed as:
$\mathbf{X}_{\text{resampled}}^{(k)}[t^{\prime}]=\mathbf{X}^{(k)}[t]+\left(\mathbf{X}^{(k)}[t+1]-\mathbf{X}^{(k)}[t]\right)\cdot(t^{\prime}-t)$
This ensures consistent temporal resolution across datasets. We use $\Delta t\approx 0.333$ seconds as a compromise to retain high- and low-frequency dynamics.
Importantly, the resampling step is the last preprocessing step necessary to create the neural activity data matrix for each worm:
$\mathbf{X}^{(k)}\leftarrow\mathbf{X}_{\text{resampled}}^{(k)}$
6. Masking and Subsetting. We also create a binary mask for each worm, $\mathbf{M}^{(k)}\in\{0,1\}^{D}$, where $D=300$ represents the total number of canonical neurons in C. elegans. Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled and measured for worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not measured.
The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$ enabling in its potential use in downstream analyses that require filtering out unlabeled neurons. Therefore, the pair $(\mathbf{M}^{(k)},\mathbf{X}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
In the next section, we describe the second half the preprocessing pipeline, which unifies many measurements of the of C. elegans into a standardized format.
Figure 1: Processed Neural Activity Data Structure. Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C. elegans. The calcium signals are normalized, smoothed, and resampled to a common time interval ($\Delta t$). A binary mask is generated to indicate which neurons were labeled for each worm. The mask does not modify the data but is retained as metadata to aid future analyses (e.g., selecting only labeled neurons).
Algorithm 1 Neural Data Preprocessing Pipeline
[…]
for each dataset $\mathcal{D}_{\mathrm{raw,source}}$ in $\mathcal{D}_{\mathrm{raw,all\_sources}}$ do
     for each worm $k$ in $\mathcal{D}_{i}$ do
         Extract calcium fluorescence traces from source files: $\mathbf{X}^{(k)}_{\text{raw}}$
         Normalize traces: $\mathbf{X}^{(k)}_{\text{norm}}\leftarrow\text{normalize}(\mathbf{X}^{(k)}_{\text{raw}})$
         (Optional) Smooth traces: $\mathbf{X}^{(k)}_{\text{smooth}}\leftarrow\text{smooth}(\mathbf{X}^{(k)}_{\text{norm}},\alpha)$
         Resample traces: $\mathbf{X}^{(k)}_{\text{resample}}\leftarrow\text{resample}(\mathbf{X}^{(k)}_{\text{smooth}},\Delta t)$
         Processed neural data: $\mathbf{X}^{(k)}\leftarrow\mathbf{X}^{(k)}_{\text{resample}}$
         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\mathbf{X}^{(k)}_{j}\mid\mathbf{X}^{(k)}_{j}\neq\emptyset,\,j\in[1,D]\}$
         Store processed neural data and labeled neuron mask: $\mathbf{X}^{(k)}$ and $\mathbf{M}^{(k)}$
     end for
     Save the processed dataset $\mathcal{D}_{\mathrm{processed,source}}:=\{\left(\mathbf{X}^{(k)},\mathbf{M}^{(k)}\right)\}_{k=1}^{N}$
end for
Combine the processed datasets
2.3 Connectome Graph Data
We processed multiple open-source C. elegans connectome datasets that capture the full neural wiring diagram, including both chemical synapses and gap junctions. Table 5 summarizes these datasets, which vary in annotation precision and recording techniques.
Table 3: Open-Source Connectome Datasets Metadata.Metadata for open-source C. elegans connectome datasets. Each dataset was standardized into the graph_tensor format, listing the corresponding graph tensor file and the total number of edges (synaptic and electrical connections) per dataset.
Source Dataset Name Graph Tensor File Number of Edges
Cook2019 [3] graph_tensors_cook2019.pt 46837
White1986 (whole) [24] graph_tensors_white1986_whole.pt 46603
White1986 (n2u) [24] graph_tensors_white1986_n2u.pt 45808
White1986 (jsh) [24] graph_tensors_white1986_jsh.pt 45759
White1986 (jse) [24] graph_tensors_white1986_jse.pt 44901
Witvliet2020 (7) [25] graph_tensors_witvliet2020_7.pt 46057
Witvliet2020 (8) [25] graph_tensors_witvliet2020_8.pt 46107
OpenWorm2023 [21] graph_tensors_openworm.pt 47683
Randi2023 [14] graph_tensors_funconn.pt 50521
Chklovskii2011 [19] graph_tensors_chklovskii.pt 46655
The connectome data is represented as a directed graph where neurons are nodes and edges represent synaptic or gap junction connections.
Figure 2: Graphical Representation of the Connectome Data. Graph-based structure used for the C. elegans connectome data. Neurons are represented as nodes, with connections (edges) classified as either chemical synapses or gap junctions. The neuron features, such as positional coordinates, neuron class, and type, are stored as node attributes, while edge attributes capture synaptic weights and connection types. This graph-based structure enables connectivity-informed modeling of neural activity. The differently colored nodes arbitrarily represent the presence of different types of neurons within the C. elegans nervous system (e.g. motor, inter & sensory neurons). The tables correspond to the bolded edge and node. The attr(e_1) feature for a single edge is the tuple of chemical synapse and gap junction weights.
2.3.1 Graph Tensor Data Processing
The data processing is handled by the ConnectomeBasePreprocessor class, which standardizes the neuron indices, processes edge attributes, and constructs the final graph tensors.
Table 4: Attributes in Graph Tensor Format. Attributes and their descriptions in the standardized graph tensor format. Includes neuron identities, edge connectivity types, synaptic weights, and spatial coordinates.
Symbol Description
$k$ Number of labeled neurons in a dataset
$l$ Number of labeled edges in a dataset
$N_{\text{pre}}^{k}$ Enumerated neuron identities for presynaptic neurons
$N_{\text{post}}^{k}$ Enumerated neuron identities for postsynaptic neurons
$T^{k}$ Connection type binary values (gap junction vs. chemical synapse)
$E$ Presynaptic neuron, postsynaptic neuron pairs which form edges
$G^{l}$ Gap junction weights for $l$ edges
$C^{l}$ Chemical synapse weights for $l$ edges
$x,y,z$ Cartesian coordinate $x,y,z$ position values for each neuron
1. Column Data Format.
Within each source connectome, presynaptic and postsynaptic neurons were specified along with chemical synapse and gap junction weight values. In the majority of the source connectome files, the format tabular for $k$ neurons with a column for the presynaptic neuron identity denoted by $N_{\text{pre}}^{k}$, a column for postsynaptic neuron identity denoted by $N_{\text{post}}^{k}$, connection type (gap junction vs. chemical synapse) denoted by $T^{k}$ with binary value elements (0 or 1), and the number of synapses, $N_{\text{weight}}^{k}$.
Table 5: Column Data Format for Connectome Dataset. Example of tabular data structure used in connectome datasets. Columns describe the origin and target neurons, connection type (gap junction or chemical synapse), number of connections, and associated neurotransmitters. The example shown is the first and last few rows of the OpenWorm [21]
connectome source file. Origin Target Type Number of Connections Neurotransmitter ADAL ADAR GapJunction 1 Generic_GJ ADAL ADFL GapJunction 1 Generic_GJ ADAL AIBL Send 1 Glutamate ADAL AIBR Send 2 Glutamate ADAL ASHL GapJunction 1 Generic_GJ VC1 VC2 Send 3 Serotonin_Acetylcholine VC1 VC2 GapJunction 6 Generic_GJ VC1 VC3 Send 1 Serotonin_Acetylcholine VC1 VC3 GapJunction 2 Generic_GJ VC1 VD1 GapJunction 1 Generic_GJ
2. Matrix Connectome Data Format.
The exception to the tabular connectome source dataset format with distinct columns for presynaptic and postsynaptic neuron identities was the functional connectome from [14] and the connectome from [3]. These connectomes were in $N\times N$ matrix format where $N$ is the number of neurons. Neurons were arranged alphabetically. The $(i,j)$ element within the matrix represents the synaptic strength proportional to elicited response of surrounding stimulated neurons for the functional connectome from [14]. Similarly, the $(i,j)$ element within the matrix represents the number of synapses between neurons $i$ and $j$ for the connectome from [3]
3. Extraction.
Each connectome has a custom class created which inherits from the parent ConnectomeBasePreprocessor class. This was due to each connectome source being formatted differently from each other requiring custom loading and preprocessing from different file types in different formats.
Within each custom subclass, we implemented a preprocess function which was designed to open the source dataset respective file type and extract the relevant data to construct the final graph tensor object. Each preprocess function ends with the graph tensor object being saved in tensor file format via the save_graph_tensor() function call. Depending on the structure of the source dataset files, each custom preprocess function contained code responsible for opening and iterating through the different blocks of the file to extract relevant connectivity and edge weights data.
We used the torch_geometric library to instantiate a Data graph object which standardized the data structure used by the various connectome sources.
4. Graph Tensor Formatting
For each source connectome dataset, we have a list of neurons $N$ and $l$ edges $E\in\{N_{pre},N_{post}\}^{l}$. The list of edges between neurons which form the connectome are saved as the edge_index attribute of the graph tensor object. Additionally, synaptic weights for both gap junctions and chemical synapses were extracted from each source connectome dataset and saved as the edge_attr attribute of the graph tensor object. The gap junction weights, denoted by $G\in R^{l}$, and chemial synapse weights, denoted by $C\in R^{l}$, were appended together to produce the edge_attr tensor attributes. These weightages provide a standardized denotion of neuron influence to better construct graph conditioned, informed, and constrained neural activity models.
5. Cross-Dataset Pairing
The source connectome datasets did not contain all of the necessary information to produce the graph tensor object which was subsequently saved as a tensor file. Specifically, information such as the neuron type used by certain chemical synapse connections between neurons were not provided with every single source connectome dataset. As a result, we obtained connectivity information from each source dataset and paired that with neuron information associated with each chemical synapse edge. Additionally, other attributes were also shared across datasets which were not present in the original source datasets such as $x,y,z$ position of neurons. This cross-dataset pairing strategy allows for a more comprehensive and standardized graph tensor object format across different source datasets.
2.3.2 Consensus Connectome Analysis
Analysis across the different connectome source datasets was performed to extract commonalities between specific source connectome chemical synapse and gap junction connections. Specifically, we created a consensus connectome that, for every neuron-to-neuron chemical synapse or gap junction connection across all source datasets, selects the maximum weight value.
3 Discussion
In this paper we present a normalized and standardized dataset for C. elegans neural activity and connectomes. We have created & consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code blocks we have provided are composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
This unified neural and connectome dataset provides a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. However, while this dataset offers numerous advantages, there are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. One notable clarification is that while C. elegans is often said to have 302 neurons, we refer to it as having 300 neurons in this work because two of these cells, CANL and CANR, are now classified as “end organs” rather than neurons. This shift in classification is supported by the work of Stefanakis et al. [17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
One of the primary limitations arises from the use of calcium fluorescence imaging as the method for measuring neural activity. Calcium imaging, while valuable for capturing broad patterns of activity across large populations of neurons, is an indirect measure of neural activity and introduces a low-pass filtering effect [6]. Calcium signals lag behind the actual electrical activity due to the slower dynamics of calcium ion concentration changes compared to the fast voltage changes associated with action potentials. This delay can be particularly problematic for studies investigating fast synaptic interactions, as important rapid neural dynamics may be lost. Thus, while calcium imaging is useful for measuring slow, global brain states, it cannot fully capture the complexity of rapid information processing that is critical for certain neural circuits [23, 20]. This highlights the need to complement calcium imaging with other techniques, such as voltage-sensitive indicators, to gain a more comprehensive understanding of neural activity.
Additionally, while the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and potentially across the lifespan of a single individual. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [3, 14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity. This limitation means that the relationship between structure and function is inferred across different animals, potentially missing individual-specific relationships between connectivity and function, despite phenotypic matching. The generation of the C. elegans connectome, as described by White et al. [24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons due to neuromodulator & neuropeptide release which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [14].
We opted not to apply smoothing to the neural activity traces for two principal reasons. First, it is likely that the original data creators have already undertaken a degree of preprocessing on their raw microscopy or imaging data, which often includes smoothing techniques to enhance signal quality. Introducing additional smoothing could risk over-processing the data, thereby obscuring critical neural dynamics. Second, the calcium fluorescence signals utilized in our analysis are inherently low-pass filtered representations of the underlying neural activity, a characteristic of the imaging modality itself. This filtering effectively diminishes high-frequency components, which are vital for capturing rapid neural events. By refraining from further smoothing, we aim to preserve the integrity and richness of the original data, allowing for a more nuanced understanding of the biological phenomena under investigation.
Resampling was chosen as a strategy to make the data comparable across experiments with different temporal resolutions. We resampled the neural activity traces to a common time step of approximately $\Delta t\approx 0.333$ seconds, which was chosen as a balance between temporal resolution and minimizing data loss due to downsampling from experiments with higher frame rates. However, this decision introduces its own limitations. There was a tradeoff between attenuating high-frequency information from datasets with higher sampling rates versus incorrectly hallucinating data from datasets with lower sampling frequencies. For experiments that originally had much higher temporal resolution, downsampling risks losing potentially valuable high-frequency dynamics. Conversely, for lower-resolution datasets, resampling can artificially inflate the temporal resolution without actually adding more meaningful information [11]. This was a necessary compromise to allow for cross-study comparisons but should be considered when interpreting the results.
On the computational side, while this dataset serves as a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using this dataset for machine learning models. The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections. Moreover, the neural dynamics data, affected by the limitations of calcium imaging, might bias models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions. Researchers should be mindful of these constraints when developing models and consider integrating additional datasets (e.g., voltage imaging) to build more comprehensive models [8].
Despite these limitations, the C. elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics. Graph neural networks (GNNs), for instance, are well-suited to handle the structural connectivity data, while recurrent neural networks (RNNs) and other time-series models can be applied to the neural activity data to capture temporal dependencies [9]. These models could inform the development of foundation models that generalize to larger and more complex systems. Moreover, the dataset’s open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
In conclusion, while this dataset offers significant opportunities for advancing our understanding of neural systems, its limitations should be carefully considered in both biological and computational contexts. Calcium imaging, structural connectomics, and the standardization choices we made each come with trade-offs that impact how the data can be used. As neural imaging technologies continue to improve and new datasets are integrated, we expect that this dataset will serve as a foundation for building more complete models of neural function, bridging the gap between simple nervous systems like C. elegans and the more complex neural architectures found in larger organisms [13, 12].
4 Conclusion
We present a standardization protocol implemented in a code repository for the preprocessing and integration of 11 C. elegans calcium fluorescence datasets and 3 primary connectome sources [24, 3, 25]. The code repository contains preprocessing classes that load and preprocess the different source datasets, handling both calcium fluorescence and connectome data. Each dataset has a class implementation that inherits from a respective parent class, which has been adapted to accommodate varying raw data formats. Once loaded, the calcium fluorescence data is further masked, normalized, resampled, and smoothed.
This unified dataset offers an accessible and standardized resource, streamlining access to data recorded under diverse experimental conditions and formats. The datasets are available for public access on Hugging Face: (1) neural data, and (2) connectome data. Additionally, the full code repository containing the preprocessing code can be accessed on GitHub at worm-graph.
Such standardization facilitates the development of models for the entire C. elegans nervous system, allowing researchers to study the interplay between structure, via the connectome, and function, via calcium fluorescence data. Through this integration, we hope to provide further insights into how neural function emerges from structural connectivity and how these dynamics might generalize to more complex nervous systems.
The open-source nature of this resource aims to further contribute to both neuroscience and AI research, accelerating efforts to model not only C. elegans but also organisms with larger and more complex nervous systems. By releasing this dataset and accompanying code, we hope to advance the growing initiative toward neural dynamics modeling and the understanding of nervous systems across species.

### v2

@@
 While neural dynamics in the brain can often be complex and difficult to measure in larger organisms, smaller systems like C.
-elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure [24, 3].
+elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure.
 C.
-elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [25].
+elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [24, 3, 25].
 This makes it an ideal organism for studying how neural dynamics arise from connectivity and how this relationship could inform both neuroscience and AI models of neural function.
@@
 elegans provides an excellent model for whole-brain functional imaging using calcium fluorescence sensors [26].
-Calcium indicators act as proxies for voltage activity, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6].
+Calcium indicators act as proxies for local computation, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6].
 Additionally, since the full connectome of C.
-elegans has been mapped via electron microscopy [24, 25], combining these two datasets—neural dynamics and connectomic structure—allows for a deep investigation of how network connectivity influences neural function.
+elegans has been mapped via electron microscopy [24, 3, 25], combining these two datasets — neural dynamics and connectivity structure — allows for a deep investigation of how network connectivity influences neural function.
 Unification of such datasets can also contribute to the broader goal of understanding how neural activity propagates through networks, ultimately informing both the development of AI-based neural systems and providing insights into how these mechanisms generalize to more complex organisms [8].
 However, the existing datasets from calcium imaging studies of C.
-elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates [9, 10].
+elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates.
 This makes it challenging to perform cross-study comparisons or build unified models.
-To address these challenges, we have created two datasets, the first combining neural dynamics from 11 neuroimaging studies and the second combining neural connectivity from 10 connectome studies of C.
-elegans.
-This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data [5].
-The primary goal of this dataset is to facilitate the training of foundation models of neural dynamics, which can be used to study how neural function arises from structure.
-The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel [15].
-It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function [3].
+To address these challenges, we have created two datasets, the first combining neural dynamics from 11 calcium maging studies and the second combining neural connectivity from 3 connectome studies of C.
+elegans [24, 3, 25].
+This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data.
+The primary goal of this dataset is to facilitate the training of computational models of C.
+elegans neural dynamics, which can be used to study how neural function is related to structure.
+The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel.
+It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function.
 We are releasing these datasets as open-source resources on the HuggingFace platform to ensure broad accessibility to researchers in both neuroscience and machine learning.
@@
 2.1 Code Repository
-We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that supports the entire preprocessing pipeline we describe in following sections.
+We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections.
 Our preprocessing is combination of utilities (classes, functions and files) that facilitate the extraction and processing of a collection of C.
@@
 The main functions in the repository are pickle_neural_data() and preprocess_connectome().
-The former handles the extraction and processing of the raw calcium data from each of our 12 neural activity source datasets into standardized neural activity data that are saved as compressed files.
+The former handles the extraction and processing of the raw calcium data from each of our 11 neural activity source datasets into standardized neural activity data that are saved as compressed files.
 The latter processes the raw connectome data from connectome source files into compressed graph tensors that are saved as compressed in a data format amenable to downstream processing by graph neural network packages.
-There are 7 distinct connectome source files as shown in Table 5 but they are all derived from 3 primary source publications [24, 3, 25].
+There are 10 distinct connectome source files as shown in Table 3 but they are all derived from 4 primary source publications [24, 3, 25, 14].
 Therefore, our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
@@
 The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$ enabling in its potential use in downstream analyses that require filtering out unlabeled neurons.
-Therefore, the pair $(\mathbf{M}^{(k)},\mathbf{X}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
+Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
 In the next section, we describe the second half the preprocessing pipeline, which unifies many measurements of the of C. elegans into a standardized format.
@@
 Columns describe the origin and target neurons, connection type (gap junction or chemical synapse), number of connections, and associated neurotransmitters.
-The example shown is the first and last few rows of the OpenWorm [21]
-connectome source file.
+The example shown is the first and last few rows of the OpenWorm [21] connectome source file
+.
 Origin Target Type Number of Connections Neurotransmitter ADAL ADAR GapJunction 1 Generic_GJ ADAL ADFL GapJunction 1 Generic_GJ ADAL AIBL Send 1 Glutamate ADAL AIBR Send 2 Glutamate ADAL ASHL GapJunction 1 Generic_GJ VC1 VC2 Send 3 Serotonin_Acetylcholine VC1 VC2 GapJunction 6 Generic_GJ VC1 VC3 Send 1 Serotonin_Acetylcholine VC1 VC3 GapJunction 2 Generic_GJ VC1 VD1 GapJunction 1 Generic_GJ
@@
 elegans neural activity and connectomes.
-We have created & consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data.
+We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data.
 The code blocks we have provided are composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
@@
 elegans is often said to have 302 neurons, we refer to it as having 300 neurons in this work because two of these cells, CANL and CANR, are now classified as “end organs” rather than neurons.
-This shift in classification is supported by the work of Stefanakis et al.
-[17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
+This shift in classification is supported by [17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
 One of the primary limitations arises from the use of calcium fluorescence imaging as the method for measuring neural activity.
@@
 The generation of the C.
-elegans connectome, as described by White et al.
-[24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation.
+elegans connectome, as described by [24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation.
 Additionally, there is evidence for wireless connections between neurons due to neuromodulator & neuropeptide release which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [14].

### v3

@@
 There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C.
-elegans).
-This is particularly interesting as this model system provides a path to bridge the gap between structure and function, from nervous system connectivity to physiology.
-However, the many existing physiology datasets, both recording and stimulation, as well as connectome datasets, are in distinct formats, requiring extra processing steps before modeling or other analysis can commence.
-Here we present a homogenized dataset of neural activity, including during stimulation, compiled from 11 neuroimaging experiments and from 10 connectome reconstructions.
-The physiology datasets, collected under varying experimental protocols, all measure neural activity via calcium fluorescence in labeled subsets of the worm’s 300 neurons.
-Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling frequency.
-The resulting dataset includes neural recordings from approximately 900 worms and 250 uniquely labeled neurons.
-The connectome datasets, collected from electron microscopy (EM) reconstructions, all contain the entire nervous system of the worm, preprocessed into a graph of connections across the neurons.
-Using our collection of datasets is facilitated through easy data sharing on HuggingFace.
-We believe that our joint dataset of physiology and connectivity will facilitate modeling, for example in terms of recurrent neural network or transformer architectures, making it easier to check how well models generalize across animals and labs.
+elegans), as this small model system provides a path to bridge the gap between nervous system structure (connectivity) and function (physiology).
+However, existing physiology datasets, whether involving passive recording or stimulation, are in distinct formats, and connectome datasets require preprocessing before analysis can commence.
+Here we compile and homogenize datasets of neural activity and connectivity.
+Our neural activity dataset is derived from 11 C.
+elegans neuroimaging experiments, while our connectivity dataset is compiled from 9 connectome annotations based on 3 primary electron microscopy studies and 1 signal propagation study.
+Physiology datasets, collected under varying protocols, measure calcium fluorescence in labeled subsets of the worm’s 300 neurons.
+Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling rate, yielding recordings from approximately 900 worms and 250 uniquely labeled neurons.
+The connectome datasets, collected from electron microscopy reconstructions, represent the entire nervous system as a graph of connections.
+Our collection is accessible on HuggingFace, facilitating analysis of the structure-function relationship in biology using modern neural network architectures and enabling cross-lab and cross-animal comparisons.
 1 Introduction
-Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience and machine learning.
-While neural dynamics in the brain can often be complex and difficult to measure in larger organisms, smaller systems like C.
-elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure.
-C.
-elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [24, 3, 25].
-This makes it an ideal organism for studying how neural dynamics arise from connectivity and how this relationship could inform both neuroscience and AI models of neural function.
-With its transparent body and well-characterized neural architecture, C.
-elegans provides an excellent model for whole-brain functional imaging using calcium fluorescence sensors [26].
-Calcium indicators act as proxies for local computation, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6].
-Additionally, since the full connectome of C.
-elegans has been mapped via electron microscopy [24, 3, 25], combining these two datasets — neural dynamics and connectivity structure — allows for a deep investigation of how network connectivity influences neural function.
-Unification of such datasets can also contribute to the broader goal of understanding how neural activity propagates through networks, ultimately informing both the development of AI-based neural systems and providing insights into how these mechanisms generalize to more complex organisms [8].
-However, the existing datasets from calcium imaging studies of C.
-elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates.
-This makes it challenging to perform cross-study comparisons or build unified models.
-To address these challenges, we have created two datasets, the first combining neural dynamics from 11 calcium maging studies and the second combining neural connectivity from 3 connectome studies of C.
-elegans [24, 3, 25].
-This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data.
-The primary goal of this dataset is to facilitate the training of computational models of C.
-elegans neural dynamics, which can be used to study how neural function is related to structure.
-The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel.
-It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function.
-We are releasing these datasets as open-source resources on the HuggingFace platform to ensure broad accessibility to researchers in both neuroscience and machine learning.
-By unifying these datasets and making them accessible, we hope to encourage the development of computational models that can capture the structure-function relationships in small nervous systems.
+Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience.
+Recent advances in computational modeling and data acquisition have reinvigorated interest in the neural system of the nematode Caenorhabditis elegans (C.
+elegans), a model organism uniquely suited for bridging the gap between neural connectivity and function.
+With its fully mapped synaptic connectome [35, 36], consisting of approximately 300 neurons [3], and its extensively studied neuronal dynamics [23, 37], C.
+elegans provides an ideal platform for investigating how neural structure constrains or influences physiology.
+Recent research has focused on leveraging the C.
+elegans connectome for generative models of neural development [30] and linking whole-brain activity to synaptic connectivity [4].
+These studies exemplify its potential for elucidating how neural signaling is determined or constrained by connectivity and for inspiring biologically grounded computational frameworks [26].
+Furthermore, the transparent body of C.
+elegans, combined with its well-characterized neural architecture, makes it an excellent candidate for whole-brain functional imaging using calcium fluorescence sensors [37].
+Calcium imaging enables the capture of neural dynamics at single-neuron resolution across many neurons simultaneously [6], serving as a proxy for neuronal activity.
+Although calcium fluctuations may primarily reflect local processes in neurites rather than whole-neuron computations [24], compact neurons like those in C.
+elegans likely rely on such localized signaling mechanisms, making this an informative approach.
+The combination of calcium imaging datasets with the full connectome, mapped via electron microscopy [35, 3, 36], offers an unprecedented opportunity to investigate how connectivity influences neural function.
+Unified datasets of structure and function can provide valuable insights into neural activity propagation, informing both the development of artificial intelligence (AI)-based neural simulations and the understanding of mechanisms that generalize to more complex organisms [8].
+However, calcium imaging datasets from C.
+elegans experiments vary widely in their experimental conditions, labeled neurons, and sampling rates, complicating cross-study comparisons and unified modeling efforts.
+Similarly, connectome datasets—whether physical or functional—require significant preprocessing to facilitate analysis or computational modeling.
+To address these challenges, we have created two integrated datasets.
+The first aggregates neural dynamics from 11 calcium imaging studies, and the second compiles connectome data from three electron microscopy wiring studies [35, 3, 36] and one functional signal propagation study [14].
+Together, these datasets combine structure and function, offering an invaluable resource for computational modelers aiming to construct biologically grounded neural network models.
+Our preprocessing pipeline standardizes neural activity data through normalization, resampling to a common rate, and consistent organization into data structures.
+Neural recordings include data from approximately 900 worms and 250 uniquely labeled neurons.
+The connectome data is represented as a graph with attributes for neuron positions, synaptic connections, and additional metadata, facilitating downstream analyses that integrate structure and function.
+These datasets are available as open-source resources on the HuggingFace platform, ensuring accessibility for researchers across neuroscience and machine learning.
+By unifying and releasing these datasets, we aim to catalyze the development of computational models that capture the structure-function relationships in small neural systems, enabling robust cross-lab and cross-animal comparisons.
 2 Methods
@@
 We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections.
-Our preprocessing is combination of utilities (classes, functions and files) that facilitate the extraction and processing of a collection of C.
-elegans neural and connectome data in a standardized format.
+Our preprocessing is combination of utilities (classes, functions, files and links) that facilitate the extraction and processing of a collection of C.
+elegans neural and connectome data into standardized formats.
 The main classes in the repository are NeuralBasePreprocessor and ConnectomeBasePreprocessor.
 The former manages processing of neural activity data as measured by calcium fluorescence imaging [33].
-The latter manages the processing of synaptic connectivity data as measured by anatomical counts from electron microscopy [24].
+The latter manages the processing of synaptic connectivity data as measured by anatomical counts from electron microscopy (EM) [35].
 The main functions in the repository are pickle_neural_data() and preprocess_connectome().
-The former handles the extraction and processing of the raw calcium data from each of our 11 neural activity source datasets into standardized neural activity data that are saved as compressed files.
-The latter processes the raw connectome data from connectome source files into compressed graph tensors that are saved as compressed in a data format amenable to downstream processing by graph neural network packages.
-There are 10 distinct connectome source files as shown in Table 3 but they are all derived from 4 primary source publications [24, 3, 25, 14].
+The pickle_neural_data() function extracts and processes raw calcium fluorescence data from each of our 11 neural activity source datasets, standardizing them into neural activity matrices that are saved as compressed files for efficient storage and retrieval.
+The preprocess_connectome() function processes raw connectome data from source files into graph tensors, which are formatted for compatibility with graph neural network (GNN) frameworks such as PyTorch Geometric [21].
+This preprocessing step facilitates downstream tasks like connectivity-informed modeling of neural activity.
+Although there are 10 distinct connectome source files, as detailed in Table 5, all are derived from 4 primary publications [35, 3, 36, 14], ensuring comprehensive coverage of C.
+elegans neural connectivity data.
 Therefore, our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
+Figure 1: Overview of Neural Activity Datasets.
+(A) Distribution of the number of worms in each experimental dataset.
+(B) Number of recorded neurons per worm compared to the total neuronal population of C.
+elegans.
+(C) Total duration of recorded neural activity for each dataset.
+(D) Average duration of recorded activity per worm, with one hour of calcium imaging as a benchmark.
+(E) Number of sampled time steps per worm, and (F) Sampling intervals for recorded neural activity.
+The horizontal dashed lines in (E) and (F) indicate the resampled sequence length ($L=314$) and time step ($\Delta t=0.333$ seconds) used in our preprocessing pipeline.
 2.2 Calcium Fluorescence Data
@@
 These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals.
-The number of worms and identified neurons varies across datasets as showb in Table 1).
-All animals were hermaphrodites at developmental stages from L4 to adulthood.
+The number of worms and identified neurons varies across datasets as shown in Table 1.
+All worms were hermaphrodites at developmental stages no earlier than L4 [29], with the most being in adulthood (Figure 5).
 Table 1: Calcium Fluorescence Neural Activity Datasets Metadata.
 Metadata for calcium fluorescence neural activity datasets collected from various sources.
-Each dataset includes information about the experimental conditions, the number of worms, the average number of labeled and recorded neurons, and the range of labeled neurons observed.
-Source Dataset Name Database Link Data Files Num. Worms Mean Num. Neurons (labeled, recorded) Num. labeled neurons (min, max)
-Kato2015 [9] osf.io/2395t/ WT_Stim.mat 12 (42, 127) (31, 51)
-Nichols2017 [12] osf.io/kbf38/ let.mat 44 (34, 108) (23, 43)
-Skora2018 [15] osf.io/za3gt/ WT_.mat 12 (46, 129) (39, 55)
-Kaplan2020 [9] osf.io/9nfhz/ Neuron2019_Data_ *.mat 19 (36, 114) (23, 51)
-Yemini2021 [26] zenodo.org/records/ 3906530 _Activity_OH.mat 49 (110, 125) (33, 179)
-Uzel2022 [18] osf.io/3vkxn Uzel_WT.mat 6 (50, 138) (46, 58)
-Dag2023 [13] tinyurl.com/github Dag2023 data/swf702_with_ id/*.h5 7 (100, 143) (87, 110)
-Atanas2023 [5] wormwideweb.org YYYY-MM-DD-.json/h5 42 (88, 136) (64, 115)
-Leifer2023 [14] tinyurl.com/drive Leifer2023 exported_data.tar.gz 108 (64, 68) (19, 98)
-Lin2023 [11] tinyurl.comdropbox Lin2023 run_prfrd_data.mat 577 (8, 8) (1, 22)
-Venkatachalam2024 [22] chemosensory-data.worm.world 2022_herm_.zip 22 (187, 187) (185, 189)
+The table includes the number of worms, the average number of labeled and recorded neurons, the range of labeled neurons observed, and summary experimental protocol details.
+Index Dataset Name Num. Worms Mean Num. Neurons (labeled, recorded) Num. Labeled Neurons (min, max) Experimental Protocol
+1 Kato2015 [10] 12 (42, 127) (31, 51) Wild-type; L4 stage; immobilized in a microfluidic device
+2 Nichols2017 [12] 44 (34, 108) (23, 43) Wild-type and npr-1 mutants; Lethargus/pre-lethargus stages; immobilized
+3 Skora2018 [15] 12 (46, 129) (39, 55) Wild-type; fasting/starvation conditions; immobilized
+4 Kaplan2020 [9] 19 (36, 114) (23, 51) Wild-type; L4 stage; freely moving and immobilized
+5 Yemini2021 [37] 49 (110, 125) (33, 179) Transgenic strain OH15500; adult stage; immobilized
+6 Uzel2022 [18] 6 (50, 138) (46, 58) Well-fed wild-type; adult stage; immobilized
+7 Dag2023 [13] 7 (100, 143) (87, 110) Wild-type; adult stage; freely moving with subsequent NeuroPAL immobilization
+8 Atanas2023 [5] 42 (88, 136) (64, 115) Transgenic strain NLS-GCaMP7f; adult stage; freely moving
+9 Leifer2023 [14] 103 (69, 122) (26, 102) Wild-type; adult stage; immobilized
+10 Lin2023 [11] 577 (8, 8) (1, 22) Wild-type; adult stage; immobilized
+11 Venkatachalam2024 [32] 22 (187, 187) (185, 189) Wild-type; adult stage; freely moving
+2.2.1 Overview of Neural Activity Datasets
+We analyzed neural activity datasets collected from 11 independent studies, which varied in terms of the number of worms recorded, the number of neurons labeled, the duration of recordings, and sampling intervals.
+Figure 1 provides an overview of these datasets, highlighting their variability in experimental conditions.
 2.2.2 Neural Data Processing
 The entire pipeline is implemented in the NeuralBasePreprocessor class, which handles dataset-specific preprocessing tasks, including file format loading, trace extraction, and metadata creation.
-Table 2: Notation Used in the Neural Data Processing Pipeline.
-Description of the mathematical notation used throughout the neural data preprocessing pipeline.
-Variables represent worm-specific indices, neural activity data structures, temporal properties, neuron counts, and preprocessing parameters.
-Symbol Description
-$k$ Worm index variable within a single dataset
-$\mathbf{X}^{(k)}$ Neural activity data for a single worm $k$
-$T_{k}$ Number of time points of data recorded for a worm $k$
-$D$ Number of neurons in C. elegans
-$\mathbf{\mu},\mathbf{\sigma}$ Mean and standard deviation over time of neural activity data for each neuron
-$\alpha$ Optional smoothing hyperparameter
-$\mathbf{M}^{(k)}$ Binary mask generated for each worm using labeled neurons applied to neural activity data
 The neural activity data is preprocessed using a custom pipeline (Figure 2) which follows these steps:
@@
 The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking.
+Adding a new source dataset to the pipeline is as easy as defining a new child preprocessor class that implements any loading and processing logic that needs to be customized to that particular source; for example, if the new source’s data is stored in a special file type.
+Figure 2: Processed Neural Activity Data Structure.
+Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C.
+elegans.
+The calcium signals are normalized, (optionally) smoothed, and resampled to a common time interval ($\Delta t$).
+A binary mask is generated to indicate which neurons were labeled for each worm.
 We represent the neural activity data for a worm $k$ by a matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of time points and $D=300$ is the number of neurons in C.
@@
 Neuron Trace Normalization.
-The first preprocessing step is to normalize or z-score the raw (i.e.
-from the source datasets) neural activity across the neuron dimension.
+The first preprocessing step is to normalize ($z$-score) the raw neural activity (i.e.
+data loaded from the source dataset files) across the neuron dimension.
 For each neuron trace $i$ in worm $k$, we compute:
 $\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{raw}}-\mathbf{\mu}}{\mathbf{\sigma}}$
-where $\mathbf{\mu},\mathbf{\sigma}\in\mathbb{R}^{D}$ are the mean and standard deviation over time for each neuron.
+where $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$ are vectors of the mean and standard deviation, respectively, for each neuron, computed across the temporal dimension.
 4. Smoothing. Optionally, the calcium traces can be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
@@
 elegans.
-Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled and measured for worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not measured.
-The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$ enabling in its potential use in downstream analyses that require filtering out unlabeled neurons.
+Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled in worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not labeled.
+In some of the source datasets we processed there were instances of calcium fluorescence measurements from neurons which were not identified with a label.
+For simplicity, we will not make a distinction in this work between unlabeled neurons and unmeasured neurons and may use the two terms interchangeably.
+The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$.
+We do not apply the mask to modify the neural data matrix (e.g.
+zeroing out measurements for unlabeled neurons) but instead retain it as a secondary data structure that can be used in downstream modeling and analysis (e.g., selecting only labeled neurons).
 Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
 In the next section, we describe the second half the preprocessing pipeline, which unifies many measurements of the of C. elegans into a standardized format.
-Figure 1: Processed Neural Activity Data Structure.
-Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C.
-elegans.
-The calcium signals are normalized, smoothed, and resampled to a common time interval ($\Delta t$).
-A binary mask is generated to indicate which neurons were labeled for each worm.
-The mask does not modify the data but is retained as metadata to aid future analyses (e.g., selecting only labeled neurons).
 Algorithm 1 Neural Data Preprocessing Pipeline
@@
          Processed neural data: $\mathbf{X}^{(k)}\leftarrow\mathbf{X}^{(k)}_{\text{resample}}$
-         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\mathbf{X}^{(k)}_{j}\mid\mathbf{X}^{(k)}_{j}\neq\emptyset,\,j\in[1,D]\}$
+         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\texttt{boolean}(\mathbf{X}^{(k)}_{j}\neq\emptyset)\mid j\in[1,D]\}$
          Store processed neural data and labeled neuron mask: $\mathbf{X}^{(k)}$ and $\mathbf{M}^{(k)}$
@@
 end for
-Combine the processed datasets
+Compile and store all processed datasets $\mathcal{D}_{\mathrm{processed,all\_sources}}$
 2.3 Connectome Graph Data
 We processed multiple open-source C.
-elegans connectome datasets that capture the full neural wiring diagram, including both chemical synapses and gap junctions.
-Table 5 summarizes these datasets, which vary in annotation precision and recording techniques.
-Table 3: Open-Source Connectome Datasets Metadata.Metadata for open-source C.
-elegans connectome datasets.
-Each dataset was standardized into the graph_tensor format, listing the corresponding graph tensor file and the total number of edges (synaptic and electrical connections) per dataset.
-Source Dataset Name Graph Tensor File Number of Edges
-Cook2019 [3] graph_tensors_cook2019.pt 46837
-White1986 (whole) [24] graph_tensors_white1986_whole.pt 46603
-White1986 (n2u) [24] graph_tensors_white1986_n2u.pt 45808
-White1986 (jsh) [24] graph_tensors_white1986_jsh.pt 45759
-White1986 (jse) [24] graph_tensors_white1986_jse.pt 44901
-Witvliet2020 (7) [25] graph_tensors_witvliet2020_7.pt 46057
-Witvliet2020 (8) [25] graph_tensors_witvliet2020_8.pt 46107
-OpenWorm2023 [21] graph_tensors_openworm.pt 47683
-Randi2023 [14] graph_tensors_funconn.pt 50521
-Chklovskii2011 [19] graph_tensors_chklovskii.pt 46655
-The connectome data is represented as a directed graph where neurons are nodes and edges represent synaptic or gap junction connections.
-Figure 2: Graphical Representation of the Connectome Data.
+elegans connectome datasets, each capturing the neural wiring diagram, including both chemical synapses and electrical synapse (also referred as gap junctions).
+Table 5 provides metadata for these datasets, which vary in annotation precision, formats, and the extent of the nervous system covered.
+2.3.1 Connectome Representation
+The connectome is represented as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
+1. Nodes ($\mathcal{V}$): Represent individual neurons.
+$\mathcal{V}:=\{\nu_{i}\}_{i=1}^{D},\quad D:=|\mathcal{V}|=300.$
+Each node $\nu_{i}$ is associated with attributes, such as:
+  • Positional Coordinates: Spatial location of the neuron in body-atlas coordinates $(x,y,z)$ [16].
+  • Neuron Class and Type: Categories such as sensory, motor, or interneuron, and neuron-specific types (e.g., ADA, IL2, VD).
+  • Neurotransmitter: Identity of neurotransmitters released by the neuron (e.g., acetylcholine, dopamine) [27].
+  • Feature Vector ($\mathbf{x}_{i}$): Optionally, numerical data such as calcium fluorescence measurements over time.
+2. Edges ($\mathcal{E}$): Represent synaptic and electrical connections between neurons.
+:=$\mathcal{E}:=\{e_{ij}\}_{k=1}^{E},\quad E:=|\mathcal{E}|.$
+Each edge $e_{ij}$ is associated with an attribute vector $\mathbf{a}_{ij}$, which can include:
+  • $c$: Chemical synapse weight.
+  • $g$: Gap junction weight.
+  • $f$: Functional connectivity strength (optional, based on optogenetic perturbations [14]).
+The neighbors of a node $\nu_{i}$ are defined as $\mathcal{N}(\nu_{i}):=\{\nu_{j}:e_{ij}\vee e_{ji}\in\mathcal{E}\}$, where $\vee$ indicates "or", ignoring directionality.
+Edges are included in $\mathcal{E}$ if any of their attributes $\mathbf{a}_{ij}$ are nonzero.
+For example, an edge may represent gap junctions ($\mathbf{a}_{ij}=[0,g]$) or chemical synapses ($\mathbf{a}_{ij}=[c,0]$).
+In cases where functional connectivity is included, the graph may become fully connected ($|\mathcal{E}|\approx D^{2}$) unless significance thresholds are applied.
+Chemical and electrical synapse weights are based on counts from EM annotations [35, 3, 36] whereas the functional connectivity is based on signal propagation [14].
+Figure 3 illustrates this graph-based representation of the nervous system, highlighting node and edge attributes.
+Table 5 summarizes the notation used in constructing and describing this graph structure.
+2.3.2 Data Extraction and Preprocessing
+The data extraction and preprocessing pipeline is implemented in the ConnectomeBasePreprocessor class, which standardizes neuron indices, processes edge attributes, and constructs the final graph tensor Data objects.
+Each source dataset is handled by a customized subclass since data extraction and preprocessing steps vary by dataset format:
+  1. 1.
+    Tabular Formats: Most datasets provide neuron pairs, connection types, and weights in a tabular structure.
+For example, the OpenWorm [22] dataset specifies presynaptic and postsynaptic neurons, connection types (chemical or gap junction), and weights.
+See Table 5 for an example.
+  2. 2.
+    Adjacency Matrix Formats: Some datasets use $D\times D$ adjacency matrices, with entries representing synapse counts ([3]) or functional connectivity values ([14]).
+  3. 3.
+    Supplementing Attributes: Some desirable attributes about the connectome (e.g., neuron positions, neurotransmitters, cell types, etc.) may not available from the source datasets.
+We obtain these from published research (e.g., neuron positions from [16], neurotransmitters and cell types from [27]) to supplement our graph structures.
+Custom classes within the preprocessing pipeline handle dataset-specific formats, ensuring consistent graph tensor outputs.
+2.3.3 Graph Tensor Format
+To facilitate downstream analyses and modeling, we standardized the connectome data into a graph tensor format compatible with graph neural network (GNN) libraries.
+Each dataset is transformed into a torch_geometric [21] Data object with the following components:
+  • edge_index: Adjacency list encoding $\mathcal{E}$.
+  • edge_attr: Attribute vectors $\mathbf{a}_{ij}=[c,g,f,\dots]$ for edges in edge_index.
+  • x: Node feature matrix $\mathbf{X}\in\mathbb{R}^{D\times L}$, initialized to zeros or populated with temporal slices of neural activity $\mathbf{X}^{k}[t:t+L]$ for some worm $k$.
+This data structure is the interface that allows integration of the neural activity data with connectivity data.
+  • kwargs: Metadata and node attributes including neuron coordinates, class, type, and neurotransmitter.
+This tensor format ensures compatibility with downstream ML models and facilitates integration of functional ($\mathbf{X}$) and structural ($\mathcal{E}$) data.
+Table 5 provides an overview of the standardized attributes.
+2.3.4 Consensus Connectome
+To address discrepancies in counts and weights of connections between datasets, we propose a consensus connectome that aggregates data across sources using majority voting.
+For each edge $e_{ij}$, the consensus weights are:
+$c_{\text{consensus}}=\max\{c_{1},c_{2},\dots,c_{n}\},\quad g_{\text{consensus}}=\max\{g_{1},g_{2},\dots,g_{n}\},$
+where $c_{i},g_{i}$ are the chemical and gap junction weights from $i$-th connectome source.
+In order to quantify the uncertainty of this connectome, we also compute a weighted standard deviation $v$ of each edge across sources and include this as an additional attribute $\mathbf{a}=[c,g,...,v]$:
+$v_{\text{consensus}}=\lambda\cdot\operatorname{std}\{c_{1},c_{2},\dots,c_{n}\}+(1-\lambda)\cdot\operatorname{std}\{g_{1},g_{2},\dots,g_{n}\}$
+where $\lambda$ is a tuneable hyperparameter determining the relative importance of the chemical versus electrical synapse weights.
+We set $\lambda=0.9$ reflecting our knowledge that annotations of gap junctions are less certain than those of chemical synapses and that gap junctions are more sparse than chemical synapses in the connectome [3].
+Having a single consensus connectome could used to model the average C.
+elegans nervous system, which has a stereotyped neurodevelopmental trajectory [35], or to quantify the inter-animal variability in network weights, which may reflect environmental influences.
+Figure 5 visualizes the consensus connectome of the 22 amphid chemosensory neurons [11].
+Figure 3: Graph-Based Representation of the Connectome.
 Graph-based structure used for the C.
 elegans connectome data.
-Neurons are represented as nodes, with connections (edges) classified as either chemical synapses or gap junctions.
-The neuron features, such as positional coordinates, neuron class, and type, are stored as node attributes, while edge attributes capture synaptic weights and connection types.
-This graph-based structure enables connectivity-informed modeling of neural activity.
-The differently colored nodes arbitrarily represent the presence of different types of neurons within the C.
-elegans nervous system (e.g.
-motor, inter & sensory neurons).
-The tables correspond to the bolded edge and node.
-The attr(e_1) feature for a single edge is the tuple of chemical synapse and gap junction weights.
-2.3.1 Graph Tensor Data Processing
-The data processing is handled by the ConnectomeBasePreprocessor class, which standardizes the neuron indices, processes edge attributes, and constructs the final graph tensors.
-Table 4: Attributes in Graph Tensor Format.
-Attributes and their descriptions in the standardized graph tensor format.
-Includes neuron identities, edge connectivity types, synaptic weights, and spatial coordinates.
-Symbol Description
-$k$ Number of labeled neurons in a dataset
-$l$ Number of labeled edges in a dataset
-$N_{\text{pre}}^{k}$ Enumerated neuron identities for presynaptic neurons
-$N_{\text{post}}^{k}$ Enumerated neuron identities for postsynaptic neurons
-$T^{k}$ Connection type binary values (gap junction vs. chemical synapse)
-$E$ Presynaptic neuron, postsynaptic neuron pairs which form edges
-$G^{l}$ Gap junction weights for $l$ edges
-$C^{l}$ Chemical synapse weights for $l$ edges
-$x,y,z$ Cartesian coordinate $x,y,z$ position values for each neuron
-1. Column Data Format.
-Within each source connectome, presynaptic and postsynaptic neurons were specified along with chemical synapse and gap junction weight values.
-In the majority of the source connectome files, the format tabular for $k$ neurons with a column for the presynaptic neuron identity denoted by $N_{\text{pre}}^{k}$, a column for postsynaptic neuron identity denoted by $N_{\text{post}}^{k}$, connection type (gap junction vs.
-chemical synapse) denoted by $T^{k}$ with binary value elements (0 or 1), and the number of synapses, $N_{\text{weight}}^{k}$.
-Table 5: Column Data Format for Connectome Dataset.
-Example of tabular data structure used in connectome datasets.
-Columns describe the origin and target neurons, connection type (gap junction or chemical synapse), number of connections, and associated neurotransmitters.
-The example shown is the first and last few rows of the OpenWorm [21] connectome source file
-.
-Origin Target Type Number of Connections Neurotransmitter ADAL ADAR GapJunction 1 Generic_GJ ADAL ADFL GapJunction 1 Generic_GJ ADAL AIBL Send 1 Glutamate ADAL AIBR Send 2 Glutamate ADAL ASHL GapJunction 1 Generic_GJ VC1 VC2 Send 3 Serotonin_Acetylcholine VC1 VC2 GapJunction 6 Generic_GJ VC1 VC3 Send 1 Serotonin_Acetylcholine VC1 VC3 GapJunction 2 Generic_GJ VC1 VD1 GapJunction 1 Generic_GJ
-2. Matrix Connectome Data Format.
-The exception to the tabular connectome source dataset format with distinct columns for presynaptic and postsynaptic neuron identities was the functional connectome from [14] and the connectome from [3].
-These connectomes were in $N\times N$ matrix format where $N$ is the number of neurons.
-Neurons were arranged alphabetically.
-The $(i,j)$ element within the matrix represents the synaptic strength proportional to elicited response of surrounding stimulated neurons for the functional connectome from [14].
-Similarly, the $(i,j)$ element within the matrix represents the number of synapses between neurons $i$ and $j$ for the connectome from [3]
-3. Extraction.
-Each connectome has a custom class created which inherits from the parent ConnectomeBasePreprocessor class.
-This was due to each connectome source being formatted differently from each other requiring custom loading and preprocessing from different file types in different formats.
-Within each custom subclass, we implemented a preprocess function which was designed to open the source dataset respective file type and extract the relevant data to construct the final graph tensor object.
-Each preprocess function ends with the graph tensor object being saved in tensor file format via the save_graph_tensor() function call.
-Depending on the structure of the source dataset files, each custom preprocess function contained code responsible for opening and iterating through the different blocks of the file to extract relevant connectivity and edge weights data.
-We used the torch_geometric library to instantiate a Data graph object which standardized the data structure used by the various connectome sources.
-4. Graph Tensor Formatting
-For each source connectome dataset, we have a list of neurons $N$ and $l$ edges $E\in\{N_{pre},N_{post}\}^{l}$.
-The list of edges between neurons which form the connectome are saved as the edge_index attribute of the graph tensor object.
-Additionally, synaptic weights for both gap junctions and chemical synapses were extracted from each source connectome dataset and saved as the edge_attr attribute of the graph tensor object.
-The gap junction weights, denoted by $G\in R^{l}$, and chemial synapse weights, denoted by $C\in R^{l}$, were appended together to produce the edge_attr tensor attributes.
-These weightages provide a standardized denotion of neuron influence to better construct graph conditioned, informed, and constrained neural activity models.
-5. Cross-Dataset Pairing
-The source connectome datasets did not contain all of the necessary information to produce the graph tensor object which was subsequently saved as a tensor file.
-Specifically, information such as the neuron type used by certain chemical synapse connections between neurons were not provided with every single source connectome dataset.
-As a result, we obtained connectivity information from each source dataset and paired that with neuron information associated with each chemical synapse edge.
-Additionally, other attributes were also shared across datasets which were not present in the original source datasets such as $x,y,z$ position of neurons.
-This cross-dataset pairing strategy allows for a more comprehensive and standardized graph tensor object format across different source datasets.
-2.3.2 Consensus Connectome Analysis
-Analysis across the different connectome source datasets was performed to extract commonalities between specific source connectome chemical synapse and gap junction connections.
-Specifically, we created a consensus connectome that, for every neuron-to-neuron chemical synapse or gap junction connection across all source datasets, selects the maximum weight value.
+Nodes represent neurons, with node attributes (e.g., position, class) and optional features.
+Directed edges capture connections, with edge attributes (e.g., chemical and gap junction weights).
 3 Discussion
-In this paper we present a normalized and standardized dataset for C.
+In this paper we present a normalized and standardized datasets for C.
 elegans neural activity and connectomes.
 We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data.
-The code blocks we have provided are composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
-This unified neural and connectome dataset provides a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system.
-However, while this dataset offers numerous advantages, there are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models.
-One notable clarification is that while C.
-elegans is often said to have 302 neurons, we refer to it as having 300 neurons in this work because two of these cells, CANL and CANR, are now classified as “end organs” rather than neurons.
-This shift in classification is supported by [17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
-One of the primary limitations arises from the use of calcium fluorescence imaging as the method for measuring neural activity.
+The code is composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
+The homogenized neural and connectome datasets provide a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system.
+While our datasets provide a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using these datasets.
+There are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models.
+We address most most of these in the next few paragraphs.
+One of the primary limitations of our neural activity dataset arises from the use of calcium fluorescence imaging as the method for measuring neural activity.
 Calcium imaging, while valuable for capturing broad patterns of activity across large populations of neurons, is an indirect measure of neural activity and introduces a low-pass filtering effect [6].
@@
 This delay can be particularly problematic for studies investigating fast synaptic interactions, as important rapid neural dynamics may be lost.
-Thus, while calcium imaging is useful for measuring slow, global brain states, it cannot fully capture the complexity of rapid information processing that is critical for certain neural circuits [23, 20].
-This highlights the need to complement calcium imaging with other techniques, such as voltage-sensitive indicators, to gain a more comprehensive understanding of neural activity.
-Additionally, while the connectome of C.
+Thus, while calcium imaging is useful for measuring slow, global brain states, it cannot fully capture the complexity of rapid information processing that may be critical in certain neural circuits [33, 20].
+This highlights the need to complement calcium-based measurements with other techniques, such as those utilizing voltage-sensitive indicators, to gain a more comprehensive understanding of neural activity.
+However, calcium dynamics may be sufficient for the study of neural signaling in C.
+elegans as most of its neurons are nearly isopotential in the steady state [23].
+While the connectome of C.
 elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics.
-One major limitation is the variability in synaptic strengths between neurons, both across individuals and potentially across the lifespan of a single individual.
-The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [3, 14].
-Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity.
-This limitation means that the relationship between structure and function is inferred across different animals, potentially missing individual-specific relationships between connectivity and function, despite phenotypic matching.
-The generation of the C.
-elegans connectome, as described by [24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation.
-Additionally, there is evidence for wireless connections between neurons due to neuromodulator & neuropeptide release which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [14].
+One major limitation is the variability in synaptic strengths between neurons, both across individuals and across the lifespan of a single individual [36].
+The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [14].
+Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity measurements.
+This creates both an opportunity and a limitation.
+The opportunity arises from the fact that any robust relationship between structure and function to be discovered from this data must generalize across different age and phenotype-matched animals.
+The limitation is that this may preclude the development of individual-specific models which are desirable for understanding learning and memory, for example.
+While having the C.
+elegans connectome [35, 3] provides a comprehensive structural map of synaptic connections, understanding the relationship between connectivity and dynamics still requires careful interpretation.
+Additionally, there is evidence for wireless connections between neurons [14] due to neuromodulation by way of neuropeptides which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [1, 31].
 We opted not to apply smoothing to the neural activity traces for two principal reasons.
@@
 This was a necessary compromise to allow for cross-study comparisons but should be considered when interpreting the results.
-On the computational side, while this dataset serves as a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using this dataset for machine learning models.
+One important clarification about our work needs to made regarding our use of $300$ for the number of neurons in C.
+elegans, whereas readers may be more familiar the number $302$ from prior literature.
+We drop the bilateral pair of canal-associated neurons (CAN) from our datasets because they completely lack chemical synapses and are now widely classified as ‘end-organs’ rather than neurons [17, 3].
 The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections.
-Moreover, the neural dynamics data, affected by the limitations of calcium imaging, might bias models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions.
-Researchers should be mindful of these constraints when developing models and consider integrating additional datasets (e.g., voltage imaging) to build more comprehensive models [8].
+Moreover, the used of calcium measurements might bias neural activity models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions.
+Researchers should be mindful of these constraints when developing models and consider integrating additional multimodal datasets to build more comprehensive models [8].
 Despite these limitations, the C.
 elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics.
-Graph neural networks (GNNs), for instance, are well-suited to handle the structural connectivity data, while recurrent neural networks (RNNs) and other time-series models can be applied to the neural activity data to capture temporal dependencies [9].
+For instance, graph neural networks (GNNs) [34] are well-suited to handle the structural connectivity data, while time-series models like structured state space models [25] can be applied to the neural activity data to capture temporal dependencies.
 These models could inform the development of foundation models that generalize to larger and more complex systems.
-Moreover, the dataset’s open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
-In conclusion, while this dataset offers significant opportunities for advancing our understanding of neural systems, its limitations should be carefully considered in both biological and computational contexts.
-Calcium imaging, structural connectomics, and the standardization choices we made each come with trade-offs that impact how the data can be used.
+Moreover, the datasets’ open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
+In conclusion, while our homogenized datasets offer significant opportunities for advancing our understanding of small neural systems, its limitations should be carefully considered in both biological and computational contexts.
+Calcium imaging, noisy connectomics, and the standardization choices we made each come with trade-offs that impact how the data can be used.
 As neural imaging technologies continue to improve and new datasets are integrated, we expect that this dataset will serve as a foundation for building more complete models of neural function, bridging the gap between simple nervous systems like C.
-elegans and the more complex neural architectures found in larger organisms [13, 12].
+elegans and the more complex neural architectures found in larger organisms.
 4 Conclusion
@@
 The datasets are available for public access on Hugging Face: (1) neural data, and (2) connectome data.
-Additionally, the full code repository containing the preprocessing code can be accessed on GitHub at worm-graph.
+Additionally, the full code repository containing the preprocessing code can be accessed on GitHub.
 Such standardization facilitates the development of models for the entire C.
@@
 elegans but also organisms with larger and more complex nervous systems.
-By releasing this dataset and accompanying code, we hope to advance the growing initiative toward neural dynamics modeling and the understanding of nervous systems across species.
+By releasing this dataset and accompanying code, we hope to advance the growing initiative toward neural structure-function modeling and the discovery of principles of neural computation across species.

### v4

@@
 Here we compile and homogenize datasets of neural activity and connectivity.
-Our neural activity dataset is derived from 11 C.
+Our neural activity dataset is derived from 12 C.
 elegans neuroimaging experiments, while our connectivity dataset is compiled from 9 connectome annotations based on 3 primary electron microscopy studies and 1 signal propagation study.
@@
 To address these challenges, we have created two integrated datasets.
-The first aggregates neural dynamics from 11 calcium imaging studies, and the second compiles connectome data from three electron microscopy wiring studies [35, 3, 36] and one functional signal propagation study [14].
+The first aggregates neural dynamics from 12 calcium imaging studies, and the second compiles connectome data from three electron microscopy wiring studies [37, 4, 38] and one functional signal propagation study [15].
 Together, these datasets combine structure and function, offering an invaluable resource for computational modelers aiming to construct biologically grounded neural network models.
@@
 2.1 Code Repository
-We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections.
+We provide an open-source code repository that implements the entire preprocessing pipeline we describe in following sections.
 Our preprocessing is combination of utilities (classes, functions, files and links) that facilitate the extraction and processing of a collection of C.
@@
 The main functions in the repository are pickle_neural_data() and preprocess_connectome().
-The pickle_neural_data() function extracts and processes raw calcium fluorescence data from each of our 11 neural activity source datasets, standardizing them into neural activity matrices that are saved as compressed files for efficient storage and retrieval.
+The pickle_neural_data() function extracts and processes raw calcium fluorescence data from each of our 12 neural activity source datasets, standardizing them into neural activity matrices that are saved as compressed files for efficient storage and retrieval.
 The preprocess_connectome() function processes raw connectome data from source files into graph tensors, which are formatted for compatibility with graph neural network (GNN) frameworks such as PyTorch Geometric [22].
 This preprocessing step facilitates downstream tasks like connectivity-informed modeling of neural activity.
-Although there are 10 distinct connectome source files, as detailed in Table 5, all are derived from 4 primary publications [35, 3, 36, 14], ensuring comprehensive coverage of C.
-elegans neural connectivity data.
-Therefore, our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
+The 10 distinct connectome source files are detailed in Table 5.
+Our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
 Figure 1: Overview of Neural Activity Datasets.
@@
 (C) Total duration of recorded neural activity for each dataset.
-(D) Average duration of recorded activity per worm, with one hour of calcium imaging as a benchmark.
-(E) Number of sampled time steps per worm, and (F) Sampling intervals for recorded neural activity.
-The horizontal dashed lines in (E) and (F) indicate the resampled sequence length ($L=314$) and time step ($\Delta t=0.333$ seconds) used in our preprocessing pipeline.
+(D) Average recording duration per worm, with one hour of calcium imaging as a benchmark.
+(E) Number of resampled time steps per worm, and (F) Pre-resampled sampling intervals for recorded neural activity.
+The horizontal dashed line in (F) indicates the target resampled time step ($\Delta t=0.333$ seconds) used in our preprocessing pipeline.
 2.2 Calcium Fluorescence Data
-We processed neural activity datasets from 11 open-source studies of C.
+We processed neural activity datasets from 12 open-source studies of C.
 elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [15, 6, 19, 39, 10, 16, 13, 11, 12, 14].
 These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals.
-The number of worms and identified neurons varies across datasets as shown in Table 1.
-All worms were hermaphrodites at developmental stages no earlier than L4 [29], with the most being in adulthood (Figure 5).
+The number of worms and identified neurons varied across datasets as shown in Table 1.
+All worms were hermaphrodites at developmental stages no earlier than L4 [31], with most being in adulthood (Figure 5).
 Table 1: Calcium Fluorescence Neural Activity Datasets Metadata.
@@
 The table includes the number of worms, the average number of labeled and recorded neurons, the range of labeled neurons observed, and summary experimental protocol details.
-Index Dataset Name Num. Worms Mean Num. Neurons (labeled, recorded) Num. Labeled Neurons (min, max) Experimental Protocol
+Index Dataset Name Num. Worms
+Mean Num.
+Neurons
+(labeled, recorded)
+Num. Labeled
+Neurons
+(min, max)
+Experimental Protocol
 1 Kato2015 [11] 12 (42, 127) (31, 51) Wild-type; L4 stage; immobilized in a microfluidic device
@@
 4 Kaplan2020 [10] 19 (36, 114) (23, 51) Wild-type; L4 stage; freely moving and immobilized
-5 Yemini2021 [37] 49 (110, 125) (33, 179) Transgenic strain OH15500; adult stage; immobilized
-6 Uzel2022 [18] 6 (50, 138) (46, 58) Well-fed wild-type; adult stage; immobilized
+5 Nejatbakhsh2020 [29] 21 (173, 175) (163, 184) Transgenic (GCaMP6); L4 to adult stages; during chemotaxis/decision-making tasks; immobilized
+6 Yemini2021 [39] 49 (110, 110) (33, 179) Transgenic (OH15500); adult stage; immobilized
+7 Uzel2022 [19] 6 (50, 138) (46, 58) Wild-type; adult stage; Well-fed but immobilized
 8 Dag2023 [14] 7 (100, 143) (87, 110) Wild-type; adult stage; freely moving with subsequent NeuroPAL immobilization
-8 Atanas2023 [5] 42 (88, 136) (64, 115) Transgenic strain NLS-GCaMP7f; adult stage; freely moving
-9 Leifer2023 [14] 103 (69, 122) (26, 102) Wild-type; adult stage; immobilized
+9 Leifer2023 [15] 110 (63, 67) (13, 98) Wild-type; adult stage; immobilized
 10 Lin2023 [12] 577 (8, 8) (1, 22) Wild-type; adult stage; immobilized
-11 Venkatachalam2024 [32] 22 (187, 187) (185, 189) Wild-type; adult stage; freely moving
+11 Flavell2023 [1] 40 (88, 136) (64, 115) Transgenic (GCaMP7f, mNeptune2.5); adult stage; freely moving active and quiescent states
+12 Venkatachalam2024 [34] 22 (187, 187) (185, 189) Wild-type; adult stage; immobilized in microfluidic chips
 2.2.1 Overview of Neural Activity Datasets
-We analyzed neural activity datasets collected from 11 independent studies, which varied in terms of the number of worms recorded, the number of neurons labeled, the duration of recordings, and sampling intervals.
-Figure 1 provides an overview of these datasets, highlighting their variability in experimental conditions.
+We analyzed neural activity datasets collected from 12 independent studies, which varied in terms of the number of worms recorded, the number of neurons labeled, the duration of recordings, and sampling intervals.
+Figure 1 provides some summary statistics on the neural activity data collected from each of these datasets.
+2.2.2 Brief Note on Notation
+In this section, we define the key mathematical notation used throughout our preprocessing pipeline and neural activity data representation.
+This note aims to provide clarity and consistency for readers, particularly when interpreting symbols used in equations and methods.
+For a comprehensive description of the symbols, including those specific to other parts of the pipeline, refer to Appendix Table 5.
+We represent the neural activity data for a worm $k$ as a time-series matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of recorded time points, and $D=300$ is the total number of canonical neurons in C.
+elegans.
+This excludes the canal-associated neurons (CAN), which are not typically included in functional studies [18, 33].
+Each row of $\mathbf{X}^{(k)}$ corresponds to a snapshot of neural activity across all $D$ neurons at a single time point, while each column corresponds to the activity trace for a specific neuron over time.
+Key conventions include:
+  • Activity Snapshot: The neural activity at time $t$ is denoted $\mathbf{X}^{(k)}[t]\in\mathbb{R}^{D}$, representing the activity of all $D$ neurons at that moment.
+  • Indexing: We use 00-based indexing for time points (e.g., $t=0,1,\dots,T_{k}-1$) and $1$-based indexing for neurons (e.g., $i=1,2,\dots,D$).
+  • Vector Orientation: All vectors (e.g., $\mathbf{X}^{(k)}[t]$) are column vectors unless explicitly stated otherwise.
+This notation provides the foundation for the mathematical operations described in the neural data processing pipeline. For additional details on notation, see Appendix Table 5.
 2.2.3 Neural Data Processing
 The entire pipeline is implemented in the NeuralBasePreprocessor class, which handles dataset-specific preprocessing tasks, including file format loading, trace extraction, and metadata creation.
-The neural activity data is preprocessed using a custom pipeline (Figure 2) which follows these steps:
-1.
+The neural activity data is preprocessed using a custom pipeline (Figure 2) which follows the general outline given in Algorithm 1 and described in more detail next.
+a).
 Data Download and Extraction.
@@
 The function download_url() downloads the dataset, and extract_zip() decompresses the data into a local directory while preserving the original folder structure.
-2.
+b).
 Custom Dataset Implementations.
 Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format.
-The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking.
+The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps in order: resampling, optional smoothing, and normalization.
 Adding a new source dataset to the pipeline is as easy as defining a new child preprocessor class that implements any loading and processing logic that needs to be customized to that particular source; for example, if the new source’s data is stored in a special file type.
@@
 elegans.
-The calcium signals are normalized, (optionally) smoothed, and resampled to a common time interval ($\Delta t$).
+The calcium signals are resampled, (optionally) smoothed, and normalized to the same rate ($1/\Delta t$ Hz).
 A binary mask is generated to indicate which neurons were labeled for each worm.
-We represent the neural activity data for a worm $k$ by a matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of time points and $D=300$ is the number of neurons in C.
-elegans.
-The neural activity at any time point $t$ is denoted as $\mathbf{X}^{(k)}[t]\in\mathbb{R}^{D}$, which represents the activity snapshot across the $D$ neurons.
-3.
-Neuron Trace Normalization.
-The first preprocessing step is to normalize ($z$-score) the raw neural activity (i.e.
-data loaded from the source dataset files) across the neuron dimension.
-For each neuron trace $i$ in worm $k$, we compute:
-$\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{raw}}-\mathbf{\mu}}{\mathbf{\sigma}}$
-where $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$ are vectors of the mean and standard deviation, respectively, for each neuron, computed across the temporal dimension.
-4. Smoothing. Optionally, the calcium traces can be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
-$\mathbf{X}^{(k)}_{\text{smooth}}[t]=\alpha\mathbf{X}^{(k)}_{\text{norm}}[t]+(1-\alpha)\mathbf{X}^{(k)}_{\text{norm}}[t-1]$
-where $\alpha$ is a smoothing hyperparameter set in preprocess.yaml.
+c).
+Resampling.
+The first preprocessing step resamples the raw data to a fixed time interval $\Delta t$, using linear interpolation for datasets with different temporal resolutions.
+The new time index $t^{\prime}$ is calculated by:
+$t^{\prime}=\left\lfloor\frac{t\cdot\Delta t_{\text{original}}}{\Delta t}\right\rfloor$
+The resampled traces $\mathbf{X}_{\text{resampled}}^{(k)}$ are computed from the raw data $\mathbf{X}^{(k)}_{\text{raw}}$ as:
+$\mathbf{X}_{\text{resampled}}^{(k)}[t^{\prime}]=\mathbf{X}^{(k)}_{\text{raw}}[t]+\left(\mathbf{X}^{(k)}_{\text{raw}}[t+1]-\mathbf{X}^{(k)}_{\text{raw}}[t]\right)\cdot(t^{\prime}-t)$
+This ensures consistent temporal resolution across datasets. We use $\Delta t\approx 0.333$ seconds as a compromise to retain high- and low-frequency dynamics. After resampling, we set:
+$\mathbf{X}^{(k)}\leftarrow\mathbf{X}_{\text{resampled}}^{(k)}$
+Resampling is done as the first preprocessing step to ensure meaningful cumulative data, as we’ll discuss later on in step 5 of preprocessing, normalization.
+d). Optional Smoothing. Following resampling, the calcium traces can optionally be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
+$\mathbf{X}^{(k)}_{\text{smooth}}[t]=\alpha\mathbf{X}^{(k)}[t]+(1-\alpha)\mathbf{X}^{(k)}[t-1]$
+where $\alpha$ is a smoothing hyperparameter set in preprocess/config.py.
 For our released neural dataset, smoothing was not applied ($\alpha=1$) as prior preprocessing had already been performed.
 Calcium fluorescence data inherently reflects low frequency neural dynamics.
-Smoothing, which acts further low pass over the raw data, is not needed.
-5. Resampling. We resample the data to a fixed time interval $\Delta t$, using linear interpolation for datasets with different temporal resolutions. The new time index $t^{\prime}$ is calculated by:
-$t^{\prime}=\left\lfloor\frac{t\cdot\Delta t_{\text{original}}}{\Delta t}\right\rfloor$
-The resampled traces $\mathbf{X}_{\text{resampled}}^{(k)}$ are computed as:
-$\mathbf{X}_{\text{resampled}}^{(k)}[t^{\prime}]=\mathbf{X}^{(k)}[t]+\left(\mathbf{X}^{(k)}[t+1]-\mathbf{X}^{(k)}[t]\right)\cdot(t^{\prime}-t)$
-This ensures consistent temporal resolution across datasets. We use $\Delta t\approx 0.333$ seconds as a compromise to retain high- and low-frequency dynamics.
-Importantly, the resampling step is the last preprocessing step necessary to create the neural activity data matrix for each worm:
-$\mathbf{X}^{(k)}\leftarrow\mathbf{X}_{\text{resampled}}^{(k)}$
-6.
+Smoothing acts as a further low pass over the resampled data and may therefore be unnecessary.
+e).
+Neuron Trace Normalization.
+As the final preprocessing step, the preprocessors normalize (or $z$-score) the neural activity across the neuron dimension.
+Let $\mathbf{X}^{(k)}_{\text{input}}:=\mathbf{X}^{(k)}_{\text{smooth}}$ if smoothing was applied; otherwise $\mathbf{X}^{(k)}_{\text{input}}:=\mathbf{X}^{(k)}_{\text{resampled}}$.
+For each neuron trace $i$ in worm $k$, we compute:
+$\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{input}}-\boldsymbol{\mu}}{\boldsymbol{\sigma}},$ (1)
+where $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$ are column vectors of the mean and standard deviation, respectively, for each neuron, calculated across the temporal dimension.
+This standard normalization method uses the full time-series to compute $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$, which introduces information from future states into earlier time points, thereby violating causality.
+While this may be acceptable for some applications, respecting causality could be crucial in certain cases, such as for real-time or streaming neural activity models.
+f).
 Masking and Subsetting.
@@
 Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
-In the next section, we describe the second half the preprocessing pipeline, which unifies many measurements of the of C. elegans into a standardized format.
 Algorithm 1 Neural Data Preprocessing Pipeline
 […]
-for each dataset $\mathcal{D}_{\mathrm{raw,source}}$ in $\mathcal{D}_{\mathrm{raw,all\_sources}}$ do
-     for each worm $k$ in $\mathcal{D}_{i}$ do
+for each dataset $\mathcal{D}_{\mathrm{raw,source}~{}i}$ in $\mathcal{D}_{\mathrm{raw,all\_sources}}$ do
+     for each worm $k$ in $\mathcal{D}_{\mathrm{raw,source}i}$ do
          Extract calcium fluorescence traces from source files: $\mathbf{X}^{(k)}_{\text{raw}}$
          Normalize traces: $\mathbf{X}^{(k)}_{\text{norm}}\leftarrow\text{normalize}(\mathbf{X}^{(k)}_{\text{raw}})$
-         (Optional) Smooth traces: $\mathbf{X}^{(k)}_{\text{smooth}}\leftarrow\text{smooth}(\mathbf{X}^{(k)}_{\text{norm}},\alpha)$
-         Resample traces: $\mathbf{X}^{(k)}_{\text{resample}}\leftarrow\text{resample}(\mathbf{X}^{(k)}_{\text{smooth}},\Delta t)$
+         if Smooth traces then
+              $\mathbf{X}^{(k)}_{\text{original}}\leftarrow\text{smooth}(\mathbf{X}^{(k)}_{\text{norm}},\alpha)$
+         else
+              $\mathbf{X}^{(k)}_{\text{original}}\leftarrow\mathbf{X}^{(k)}_{\text{norm}}$
+         end if
+         Resample traces: $\mathbf{X}^{(k)}_{\text{resample}}\leftarrow\text{resample}(\mathbf{X}^{(k)}_{\text{original}},\Delta t)$
          Processed neural data: $\mathbf{X}^{(k)}\leftarrow\mathbf{X}^{(k)}_{\text{resample}}$
-         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\texttt{boolean}(\mathbf{X}^{(k)}_{j}\neq\emptyset)\mid j\in[1,D]\}$
+         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\texttt{boolean}(\mathbf{X}^{(k)}_{j}\neq\boldsymbol{\emptyset})\mid j\in[1,D]\}$
          Store processed neural data and labeled neuron mask: $\mathbf{X}^{(k)}$ and $\mathbf{M}^{(k)}$
      end for
-     Save the processed dataset $\mathcal{D}_{\mathrm{processed,source}}:=\{\left(\mathbf{X}^{(k)},\mathbf{M}^{(k)}\right)\}_{k=1}^{N}$
+     Save the processed dataset $\mathcal{D}_{\mathrm{processed,source}~{}i}:=\{\left(\mathbf{X}^{(k)},\mathbf{M}^{(k)}\right)\}_{k=1}^{N}$
 end for
@@
 We processed multiple open-source C.
-elegans connectome datasets, each capturing the neural wiring diagram, including both chemical synapses and electrical synapse (also referred as gap junctions).
+elegans connectome datasets, each capturing the neural wiring diagram, including both chemical synapses and electrical synapse (also known as gap junctions).
 Table 5 provides metadata for these datasets, which vary in annotation precision, formats, and the extent of the nervous system covered.
@@
   • Positional Coordinates: Spatial location of the neuron in body-atlas coordinates $(x,y,z)$ [17].
-  • Neuron Class and Type: Categories such as sensory, motor, or interneuron, and neuron-specific types (e.g., ADA, IL2, VD).
+  • Neuron Class and Type: Categories such as sensory, motor, or interneuron, and neuron-specific types (e.g., ADA, IL2, VD1).
   • Neurotransmitter: Identity of neurotransmitters released by the neuron (e.g., acetylcholine, dopamine) [28].
-  • Feature Vector ($\mathbf{x}_{i}$): Optionally, numerical data such as calcium fluorescence measurements over time.
+  • Feature Vector ($\mathbf{x}_{i}$): Optionally, numerical data (e.g. calcium fluorescence measurements over a fixed time window, gene expression embeddings)
 2. Edges ($\mathcal{E}$): Represent synaptic and electrical connections between neurons.
@@
 For example, an edge may represent gap junctions ($\mathbf{a}_{ij}=[0,g]$) or chemical synapses ($\mathbf{a}_{ij}=[c,0]$).
-In cases where functional connectivity is included, the graph may become fully connected ($|\mathcal{E}|\approx D^{2}$) unless significance thresholds are applied.
-Chemical and electrical synapse weights are based on counts from EM annotations [35, 3, 36] whereas the functional connectivity is based on signal propagation [14].
+In cases where functional connectivity is included, the graph may become fully connected ($|\mathcal{E}|\approx D^{2}$) unless a significance threshold is applied to filter out weak functional connections.
+Chemical and electrical synapse weights are based on counts from EM annotations [37, 4, 38] whereas the functional connectivity weights are based on signal propagation [15].
 Figure 3 illustrates this graph-based representation of the nervous system, highlighting node and edge attributes.
 Table 5 summarizes the notation used in constructing and describing this graph structure.
+Figure 3: Graph-Based Representation of the Connectome.
+Graph-based structure used for the C.
+elegans connectome data.
+Nodes represent neurons, with node attributes (e.g., position, class) and optional features.
+Directed edges capture connections, with edge attributes (e.g., chemical and gap junction weights).
 2.3.2 Data Extraction and Preprocessing
@@
   3. 3.
-    Supplementing Attributes: Some desirable attributes about the connectome (e.g., neuron positions, neurotransmitters, cell types, etc.) may not available from the source datasets.
+    Supplementing Attributes: Some desirable attributes about the connectome (e.g., neuron positions, neurotransmitters, cell types, etc.) may not be available from the source datasets.
 We obtain these from published research (e.g., neuron positions from [17], neurotransmitters and cell types from [28]) to supplement our graph structures.
@@
 2.3.3 Graph Tensor Format
-To facilitate downstream analyses and modeling, we standardized the connectome data into a graph tensor format compatible with graph neural network (GNN) libraries.
+To facilitate downstream analyses and modeling, we standardized the connectome data into a graph tensor format compatible with the PyTorch Geometric graph neural network (GNN) libraries.
 Each dataset is transformed into a torch_geometric [22] Data object with the following components:
@@
 2.3.4 Consensus Connectome
-To address discrepancies in counts and weights of connections between datasets, we propose a consensus connectome that aggregates data across sources using majority voting.
-For each edge $e_{ij}$, the consensus weights are:
-$c_{\text{consensus}}=\max\{c_{1},c_{2},\dots,c_{n}\},\quad g_{\text{consensus}}=\max\{g_{1},g_{2},\dots,g_{n}\},$
-where $c_{i},g_{i}$ are the chemical and gap junction weights from $i$-th connectome source.
-In order to quantify the uncertainty of this connectome, we also compute a weighted standard deviation $v$ of each edge across sources and include this as an additional attribute $\mathbf{a}=[c,g,...,v]$:
-$v_{\text{consensus}}=\lambda\cdot\operatorname{std}\{c_{1},c_{2},\dots,c_{n}\}+(1-\lambda)\cdot\operatorname{std}\{g_{1},g_{2},\dots,g_{n}\}$
-where $\lambda$ is a tuneable hyperparameter determining the relative importance of the chemical versus electrical synapse weights.
-We set $\lambda=0.9$ reflecting our knowledge that annotations of gap junctions are less certain than those of chemical synapses and that gap junctions are more sparse than chemical synapses in the connectome [3].
-Having a single consensus connectome could used to model the average C.
-elegans nervous system, which has a stereotyped neurodevelopmental trajectory [35], or to quantify the inter-animal variability in network weights, which may reflect environmental influences.
-Figure 5 visualizes the consensus connectome of the 22 amphid chemosensory neurons [11].
-Figure 3: Graph-Based Representation of the Connectome.
-Graph-based structure used for the C.
-elegans connectome data.
-Nodes represent neurons, with node attributes (e.g., position, class) and optional features.
-Directed edges capture connections, with edge attributes (e.g., chemical and gap junction weights).
+To address discrepancies in connection counts and weights across datasets, we construct a consensus connectome that aggregates synaptic data from multiple sources using a directional averaging strategy.
+Unlike the dataset-specific files used during preprocessing — which store information in graph network format with edge indices and attributes — the consensus connectome is a flat .csv table, not a graph.
+This tabular format is universal, self-explanatory, and easily readable by both humans and software tools.
+Each row corresponds to a directed pair of neurons, with clearly defined anatomical, connection, and metadata fields.
+This structure is especially suited for accessible downstream analysis, visualization, and integration, making it the preferred format for summarizing the overall state of connectivity across datasets.
+For each ordered pair of neurons $(i,j)$, we compute the consensus weights as follows:
+  • Gap Junctions are symmetric. Their consensus weight $g_{\text{consensus}}$ is computed as the mean of all non-zero gap weights observed in either direction across datasets:
+    $g_{\text{consensus}}=\operatorname{mean}\{g_{ij}^{(k)}\}$
+    where $g_{ij}^{(k)}$ is the gap weight from neuron $i$ to $j$ in dataset $k$. By definition, $g_{ij}^{(k)}=g_{ji}^{(k)}$.
+  • Chemical Synapses are directional.
+Their consensus weight $c_{\text{consensus}}$ is the mean of all non-zero chemical weights from neuron $i$ to $j$, excluding values from the functional connectivity dataset:
+    $c_{\text{consensus}}=\operatorname{mean}\{c_{ij}^{(k)}\}_{k\neq\text{funconn}}.$
+  • Functional Connectivity is reported directly from the dedicated funconn dataset, if available. These values are not averaged and apply only in the directed case $i\rightarrow j$.
+To quantify the uncertainty in gap and chemical connection estimates, we compute a weighted standard deviation $v$ that reflects both gap and chemical synapse variability:
+$v_{\text{consensus}}=\lambda\cdot\operatorname{std}(\{c_{ij}^{(k)}\}_{k\neq\text{funconn}})+(1-\lambda)\cdot\operatorname{std}(\{g_{ij}^{(k)}\}_{k}),$
+where $\lambda=0.9$ is a subjective value chosen to weight the uncertainty from chemical synapses more due to their relative abundance compared to gap junctions.
+This consensus connectome includes all neuron pairs — including self-loops — and supports multiple output formats (e.g., trimmed vs.
+full, with or without functional weights).
+It provides a representative model of the stereotyped C.
+elegans nervous system [37], while capturing cross-dataset variability that may reflect biological and technical noise.
+Figure 5 visualizes the consensus connectivity among the 22 amphid chemosensory neurons [12].
 3 Discussion
-In this paper we present a normalized and standardized datasets for C.
+In this paper we present standardized datasets for C.
 elegans neural activity and connectomes.
@@
 While our datasets provide a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using these datasets.
-There are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models.
-We address most most of these in the next few paragraphs.
+There are several important limitations and methodological choices that must be considered when interpreting the data and applying it in bio-computational models.
+We address most of these next.
 One of the primary limitations of our neural activity dataset arises from the use of calcium fluorescence imaging as the method for measuring neural activity.
@@
 elegans connectome [37, 4] provides a comprehensive structural map of synaptic connections, understanding the relationship between connectivity and dynamics still requires careful interpretation.
-Additionally, there is evidence for wireless connections between neurons [14] due to neuromodulation by way of neuropeptides which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [1, 31].
+Additionally, there is evidence for wireless connections between neurons [15] due to neuromodulation by way of neuropeptides which can bind to nearby and distant neuron GPCR sites triggering secondary messenger pathways [2, 33].
 We opted not to apply smoothing to the neural activity traces for two principal reasons.
-First, it is likely that the original data creators have already undertaken a degree of preprocessing on their raw microscopy or imaging data, which often includes smoothing techniques to enhance signal quality.
-Introducing additional smoothing could risk over-processing the data, thereby obscuring critical neural dynamics.
+First, it is likely that the original data creators have already performed a significant amount preprocessing on the signals extracted from their raw microscopy images, which often includes smoothing techniques to enhance signal quality.
+Introducing additional smoothing could risk over-processing the data, obscuring critical neural dynamics.
 Second, the calcium fluorescence signals utilized in our analysis are inherently low-pass filtered representations of the underlying neural activity, a characteristic of the imaging modality itself.
@@
 This was a necessary compromise to allow for cross-study comparisons but should be considered when interpreting the results.
-One important clarification about our work needs to made regarding our use of $300$ for the number of neurons in C.
+One important clarification about our work needs to be made regarding our use of $300$ for the number of neurons in C.
 elegans, whereas readers may be more familiar the number $302$ from prior literature.
@@
 The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections.
-Moreover, the used of calcium measurements might bias neural activity models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions.
+Moreover, the use of calcium measurements might bias neural activity models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions.
 Researchers should be mindful of these constraints when developing models and consider integrating additional multimodal datasets to build more comprehensive models [9].
@@
 elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics.
-For instance, graph neural networks (GNNs) [34] are well-suited to handle the structural connectivity data, while time-series models like structured state space models [25] can be applied to the neural activity data to capture temporal dependencies.
+For instance, graph neural networks (GNNs) [36] are well-suited to handle the structural connectivity data, while time-series models like structured state space models (SSMs) [26] can be applied to the neural activity data to capture temporal dependencies.
 These models could inform the development of foundation models that generalize to larger and more complex systems.
@@
 4 Conclusion
-We present a standardization protocol implemented in a code repository for the preprocessing and integration of 11 C.
-elegans calcium fluorescence datasets and 3 primary connectome sources [35, 3, 36].
+We present a standardization protocol implemented in a code repository for the preprocessing and integration of 12 C.
+elegans calcium fluorescence datasets and 4 primary connectome publications [37, 4, 38, 15].
 The code repository contains preprocessing classes that load and preprocess the different source datasets, handling both calcium fluorescence and connectome data.